# Massive  Index  Modulation  with  Combinatorics-Free Detection  for  Integrated  Sensing  and  Communications

### v1

@@ IV Proposed Low-Complexity Detection for ISAC System via Massive Index Modulation
$=\frac{\sum_{\mathbf{e}_{t}\in\mathcal{E}}\!\big{(}\mathbf{e}_{t}\cdot\mathbb{P}_{\mathsf{e_{k_{p}}\!\!}}(\mathbf{e}_{k_{p}}\!\!=\mathbf{e}_{t})\cdot\mathbb{P}(b_{p:n}^{\mathrm{R}}|\mathbf{e}_{k_{p}})\big{)}}{\sum_{\mathbf{e}^{\prime}_{t}\in\mathcal{E}}\!\big{(}\mathbb{P}_{\mathsf{e^{\prime}_{k_{p}}\!\!}}(\mathbf{e}^{\prime}_{k_{p}}\!\!=\mathbf{e}^{\prime}_{t})\cdot\mathbb{P}(b_{p:n}^{\mathrm{R}}|\mathbf{e}^{\prime}_{k_{p}})\big{)}},\vspace{-0.5ex}$
@@
$\hat{\mathbf{e}}_{k^{\mathrm{R}}_{p}:n}\!=\!\frac{\mathbf{r}_{p}\!\odot\!{\mathbf{z}^{\mathrm{R}}_{p:n}}}{\mathbf{r}_{p}^{\mathsf{T}}\cdot{\mathbf{z}^{\mathrm{R}}_{p:n}}}\in\mathbb{R}^{N_{T}\times 1},\vspace{-0.5ex}$ (37)
@@
${\mathrm{diag}}\!\left({\;\!\mathbf{\Lambda}_{p:n}^{\mathrm{R}}}\right)=|s_{p}^{\mathrm{R}}|^{2}\!\sum_{n^{\prime}\neq n}^{2N_{R}}\!\frac{|\mathbf{h}_{n^{\prime}}^{\mathrm{R}}|^{2}_{\!\odot}}{\nu^{\mathrm{R}}_{p:{n^{\prime}}}}\vspace{-0.5ex}$ (39)
@@
$\hat{\mathbf{e}}_{k^{\mathrm{R}}_{p}:n}^{[\tau+1]}\longleftarrow\rho\cdot\hat{\mathbf{e}}_{k^{\mathrm{R}}_{p}:n}^{[\tau]}+(1-\rho)\cdot\hat{\mathbf{e}}_{k^{\mathrm{R}}_{p}:n}^{[\tau+1]},$ (40)
@@
which is equivalent to equation (33) without the self-interference cancellation.
@@
$\bm{\eta}_{p}^{\mathrm{R}}=s^{\mathrm{R}}_{p}\!\sum_{n=1}^{2N_{R}}\frac{\bar{y}^{\mathrm{R}}_{p:n}}{\nu^{\mathrm{R}}_{p:n}}\mathbf{h}_{n}^{\mathrm{R}},$ (42)
${\mathrm{diag}}\!\left({\;\!\!\mathbf{\Lambda}_{p}^{\mathrm{R}}}\right)\!=\!|s_{p}^{\mathrm{R}}|^{2}\!\sum_{n=1}^{2N_{R}}\!\frac{|\mathbf{h}_{n}^{\mathrm{R}}|^{2}_{\!\odot}}{\nu^{\mathrm{R}}_{p:{n}}},$ (43)
@@
$\tilde{\mathbf{{e}}}_{k^{\mathrm{R}}_{p}}=\underset{\mathbf{e}_{t}\in\mathcal{E}}{\mathrm{argmax}}~{}\mathbb{P}(b_{p}^{\mathrm{R}}|\mathbf{e}_{t}),$ (44)
@@
$\hat{\mathbf{{e}}}_{k^{\mathrm{R}}_{p}}\!=\!\frac{\mathbf{r}_{p}\!\odot\!{\mathbf{z}^{\mathrm{R}}_{p}}}{\mathbf{r}_{p}^{\mathsf{T}}\cdot{\mathbf{z}^{\mathrm{R}}_{p}}},$ (45)
@@
${\mathbf{z}}^{\mathrm{R}}_{p}\!=\!{\mathrm{exp}}_{\odot}\!\big{(}\bm{\eta}^{\mathrm{R}}_{p}\!-\!\tfrac{1}{2}{\mathrm{diag}}\!\left({\mathbf{\Lambda}^{\mathrm{R}}_{p}}\right)\!\big{)},$ (46)
@@
$k^{\mathrm{R}}_{p}=\underset{t\in\mathcal{T}}{\mathrm{argmax}}(\hat{\mathrm{e}}_{k^{\mathrm{R}}_{p}(t)}),$ (47)
where $\hat{\mathrm{e}}_{k^{\mathrm{R}}_{p}(t)}$ denote the $t$-th element of the consensus soft-replica vector $\hat{\mathbf{{e}}}_{k^{\mathrm{R}}_{p}}$.
@@ V Improving the Proposed UVD-GaBP Decoder
On the other hand, an errorneous estimation may arise when a soft-replica converges to the wrong unit vector, as illustrated in Figure 4b, leading to a duplicate estimate. Due to the combinatoric selection of the indices in a GQSM index vector and of general IM schemes as described in Section II-A, no duplicate indices may exist within a single index vector.
In light of the above, two sources of such errorneous duplicate convergence are identified: a) high transmit symbol similarity, and b) lack of joint/dependent distribution information between the unit vector variables. In order to mitigate the error from the above identified sources, three modifications to the GQSM transmitter and the UVD-GaBP decoder are proposed in the following.
@@ V-A IQ-Orthogonal Constellation Design
Let us assume that multiple transmit symbols have high similarity, i.e., $|s^{\mathrm{R}}_{p}-s^{\mathrm{R}}_{p^{\prime}}|^{2}\!\approx\!0$. Such behavior is illustrated by observing Fig. 4b and considering $s^{\mathrm{R}}_{p}\approx s^{\mathrm{R}}_{p^{\prime}}$, where the initialized effective symbols $s^{\mathrm{R}}_{p}\cdot\hat{\mathbf{e}}^{\mathrm{R}}_{k_{p}}$ and $s^{\mathrm{R}}_{p^{\prime}}\cdot\hat{\mathbf{e}}^{\mathrm{R}}_{k_{p^{\prime}}}$ are consequently very similar in value. Therefore, the effective symbols appearing in the core MP steps of the UVD-GaBP decoder in equations (28)-(35), may converge to either of the true symbols $s^{\mathrm{R}}_{p}\cdot{\mathbf{e}}^{\mathrm{R}}_{k_{p}}$ or $s^{\mathrm{R}}_{p^{\prime}}\cdot{\mathbf{e}}^{\mathrm{R}}_{k_{p^{\prime}}}$ and potentially produce an errorneous duplicate estimate.
@@
In order to avoid such duplicate symbols in the IQ domain, we propose a simple rotation scheme to $\mathcal{S}^{\mathrm{c}}$, which leads to non-duplicate PAM symbols in both IQ domains. Furthermore, an optimal rotation angle $\theta^{*}$ is obtained to maximize the minimum Euclidean distance between the symbols of the PAM constellation. There exist an optimal rotation angle $\theta^{*}$ for each of the quadrants, but we only consider the rotation angle within the first quadrant, i.e., $\theta\in(0,\frac{\pi}{2})$, which are shown in Table I for some sizes of the constellation.
$M$ 4 16 32 64 128 256
$\theta^{*}$ 0.464 0.245 0.165 0.124 0.082 0.062
The optimization problem is described by
$D_{\mathrm{min}\!}(\mathcal{S})\triangleq\!\!\underset{s_{m\vphantom{{}^{\prime}}}\!,s_{m^{\prime}}\in\mathcal{S}}{\mathrm{min}}\,\sqrt{|s_{m\vphantom{{}^{\prime}}}\!-s_{m^{\prime}}|^{2}},~{}\forall m\neq m^{\prime}.$ (49)
@@ V-B Conditional PDF-based Denoiser
$\check{p}^{\mathrm{R}}_{n},\check{t}^{\mathrm{R}}_{n}=\underset{p,t}{\mathrm{argmax}}\;\mathbb{P}(b_{p:n}^{\mathrm{R}}|\mathbf{e}_{t}),\vspace{-0.5ex}$ (50)
@@
The conditional PMFs does not exhibit a closed-form expression unlike equation (23) for the independent PMFs, and therefore must be obtained via numerical evaluation. However, as the conditional PMFs are fixed for a given codebook, i.e., system parameters $N_{T}$ and $P$, all values of the conditional PMF can be pre-computed.
@@ V-C Greedy Successive Interference Cancellation (IC)
$\check{p}^{\mathrm{R}},\check{t}^{\mathrm{R}}=\underset{p,t}{\mathrm{argmax}}\;\mathbb{P}(b_{p}^{\mathrm{R}}|\mathbf{e}_{t}),\vspace{-0.5ex}$ (52)
@@
In hand of a converged index of the most confident unit vector variable, IC can performed on the received signals corresponding to the selected indices as
$\check{y}^{\mathrm{R}}_{n}\triangleq y_{n}-(\mathbf{h}_{n}^{\mathrm{R}})^{\!\mathsf{T}}(s^{\mathrm{R}}_{\check{p}^{\mathrm{R}}_{n}}\!\cdot\!\mathbf{e}_{\check{t}^{\mathrm{R}}_{n}}),\vspace{-0.5ex}$ (53)
and the corresponding symbol and unit vector variables are nullified from the respective soft-replica, to be repeated for $P$ successive IC iterations.
@@
$\tilde{\mathbf{k}}^{\mathrm{R}}=\mathrm{sort}([{\check{t}^{\mathrm{R}}_{[1]}},\cdots,{\check{t}^{\mathrm{R}}_{[P]}}]),$ (54)
@@
The complete procedure, incorporating the proposed receiver-side enhancements of Sections V-C and V-B, is provided as a pseudocode in Algorithm 2, and as a schematic diagram in Figure 6. Notice that for $P=1$, the proposed enhancement is not necessary as there is only a single unit vector to evaluate, and the conditional PMF-based denoiser is not applicable as there is no other unit vector to condition the extrinsic beliefs. In addition, the motivation of such algorithmic enhancements is not necessary for $P=1$, as the duplicate convergence error cannot occur with a single unit vector variable.
@@ VI-A BER Performance
To analyze the performance of the proposed UVD-GaBP algorithm and the enhancement modules in Sec. V, we perform simulations for different system sizes characterized by $N_{T}$ transmit antennas, $N_{R}$ receive antennas, and $P$ pilot symbols from the $M$-ary constellation, considering the three cases specified in Table II. The cases are selected to capture and highlight the improvements introduced by each of the three proposed enhancement modules, namely: Algorithm 1 as the baseline UVD-GaBP decoder; Algorithm 2 with conditional denoiser but without successive IC; and finally the full Algorithm 2 with both the conditional denoiser and successive IC. Optimized constellations are employed, however, in all cases considered.
@@
Our first set of results, depicted in Figure 8, compares the BER performances of the proposed UVD-GaBP decoder for all cases in Table II and the corresponding MFB, in MIMO systems of size $32\times 32$ employing various numbers of symbols $P$. The BER plots, shown against the energy-per-bit-to-noise-power-spectral-density ratio ($E_{b}/N_{0}$), indicate that the duplicate convergence of unit vectors, identified in Section V as a critical potential source of errors, does not occur for $P=1$ but can be substantial for $P>1$. The results also show, however, that the duplicate convergence behavior is significantly mitigated by the improvements described in Secs. V-B and V-C.
The BER floor behavior observed for $P>1$ is, however, not only a consequence of duplicate convergence of unit vectors, but also a function of system size. To clarify, larger systems can sustain larger $P$ values, without exhibiting error floors. In order to demonstrate that, we offer our next set of results, given in Figure 8, which show the performances achieved by Algorithm 2 in systems of size $64\times 64$ and $96\times 96$. We remark that simulation results for GQSM systems of such sizes have never been shown before, precisely due to the prohibitive complexity of SotA GQSM detection methods [46, 39].
All in all, given that error floors do not affect large systems, and since the decoding complexity of such large systems is tamed by the proposed method one can say that the proposed UVD-GaBP offers a systematic mechanism to employ massive MIMO in order to achieve high-performance ISAC functionalities with controlled complexity.
Notice that in Figure 8 we compare the performance of the proposed piloted GQSM scheme against a conventional multiplexed (MUX) MIMO-ISAC scheme, in which all antennas are active, but used to transmit symbols and pilots simultaneously, for the proposes of communications and sensing, respectively.
In order to enable a fair comparison between the two (proposed and conventional) approaches, we set the total power and transmission rate $B_{\mathrm{Tr}}$ (in bits/s/Hz) of both systems to identical values, and employ the well-known linear GaBP [70, 68, 71] in the detection of the MUX scheme. Notice, furthermore, that the proposed UVD-GaBP-based MIMO-ISAC scheme is interference-free, unlike the conventional MUX-based approach, in which decoding errors made on communication symbol estimates may deteriorate sensing performance [72, 63], while channel estimation errors made over pilots may impact on BER performance [73, 74], respectively. To the advantage of the conventional scheme, however, no interference is considered in evaluating the performance of the MUX-based system in Figure 8.
It can be seen that, piloted GQSM in large systems, enabled by the proposed UVD-GaBP detector, outperforms a conventional MIMO-MUX system in which resources are split between communcations and sensing functionalities.

### v2

@@ IV Proposed Low-Complexity Detector
$=\frac{\sum_{\mathbf{e}_{t}\in\mathcal{E}}\!\big{(}\mathbf{e}_{t}\cdot\mathbb{P}_{\mathsf{e_{k_{p}}\!\!}}(\mathbf{e}_{k_{p}}\!\!=\mathbf{e}_{t})\cdot\mathbb{P}(b_{p:n}^{\mathrm{R}}|\mathbf{e}_{k_{p}})\big{)}}{\sum_{\mathbf{e}^{\prime}_{t}\in\mathcal{E}}\!\big{(}\mathbb{P}_{\mathsf{e^{\prime}_{k_{p}}\!\!}}(\mathbf{e}^{\prime}_{k_{p}}\!\!=\mathbf{e}^{\prime}_{t})\cdot\mathbb{P}(b_{p:n}^{\mathrm{R}}|\mathbf{e}^{\prime}_{k_{p}})\big{)}},\vspace{-1.5ex}$
@@
$\hat{\mathbf{e}}_{k^{\mathrm{R}}_{p}:n}\!=\!\frac{\mathbf{r}_{p}\!\odot\!{\mathbf{z}^{\mathrm{R}}_{p:n}}}{\mathbf{r}_{p}^{\mathsf{T}}\cdot{\mathbf{z}^{\mathrm{R}}_{p:n}}}\in\mathbb{R}^{N_{T}\times 1},\vspace{-1ex}$ (37)
@@
${\mathrm{diag}}\!\left({\;\!\mathbf{\Lambda}_{p:n}^{\mathrm{R}}}\right)=|s_{p}^{\mathrm{R}}|^{2}\!\sum_{n^{\prime}\neq n}^{2N_{R}}\!\frac{|\mathbf{h}_{n^{\prime}}^{\mathrm{R}}|^{2}_{\!\odot}}{\nu^{\mathrm{R}}_{p:{n^{\prime}}}}\vspace{-1ex}$ (39)
@@
$\hat{\mathbf{e}}_{k^{\mathrm{R}}_{p}:n}^{[\tau+1]}\longleftarrow\rho\cdot\hat{\mathbf{e}}_{k^{\mathrm{R}}_{p}:n}^{[\tau]}+(1-\rho)\cdot\hat{\mathbf{e}}_{k^{\mathrm{R}}_{p}:n}^{[\tau+1]},\vspace{-0.5ex}$ (40)
@@
which is equivalent to equation (33) without the self-interference cancellation operation.
@@
$\bm{\eta}_{p}^{\mathrm{R}}=s^{\mathrm{R}}_{p}\!\sum_{n=1}^{2N_{R}}\frac{\bar{y}^{\mathrm{R}}_{p:n}}{\nu^{\mathrm{R}}_{p:n}}\mathbf{h}_{n}^{\mathrm{R}},\vspace{-0.75ex}$ (42)
${\mathrm{diag}}\!\left({\;\!\!\mathbf{\Lambda}_{p}^{\mathrm{R}}}\right)\!=\!|s_{p}^{\mathrm{R}}|^{2}\!\sum_{n=1}^{2N_{R}}\!\frac{|\mathbf{h}_{n}^{\mathrm{R}}|^{2}_{\!\odot}}{\nu^{\mathrm{R}}_{p:{n}}},\vspace{-0.25ex}$ (43)
@@
$\tilde{\mathbf{{e}}}_{k^{\mathrm{R}}_{p}}=\underset{\mathbf{e}_{t}\in\mathcal{E}}{\mathrm{argmax}}~{}\mathbb{P}(b_{p}^{\mathrm{R}}|\mathbf{e}_{t}),\vspace{-0.75ex}$ (44)
@@
$\hat{\mathbf{{e}}}_{k^{\mathrm{R}}_{p}}\!=\!\frac{\mathbf{r}_{p}\!\odot\!{\mathbf{z}^{\mathrm{R}}_{p}}}{\mathbf{r}_{p}^{\mathsf{T}}\cdot{\mathbf{z}^{\mathrm{R}}_{p}}},\vspace{-1ex}$ (45)
@@
${\mathbf{z}}^{\mathrm{R}}_{p}\!=\!{\mathrm{exp}}_{\odot}\!\big{(}\bm{\eta}^{\mathrm{R}}_{p}\!-\!\tfrac{1}{2}{\mathrm{diag}}\!\left({\mathbf{\Lambda}^{\mathrm{R}}_{p}}\right)\!\big{)},\vspace{-0.5ex}$ (46)
@@
$k^{\mathrm{R}}_{p}=\underset{t\in\mathcal{T}}{\mathrm{argmax}}(\hat{\mathrm{e}}_{k^{\mathrm{R}}_{p}(t)}),\vspace{-0.5ex}$ (47)
where $\hat{\mathrm{e}}_{k^{\mathrm{R}}_{p}(t)}$ denote the $t$-th element of $\hat{\mathbf{{e}}}_{k^{\mathrm{R}}_{p}}$.
@@ V Improving the Proposed UVD-GaBP Decoder
On the other hand, an erroneous estimation may arise when a soft-replica converges to the wrong unit vector, as illustrated in Figure 4b, leading to a duplicate estimate. Due to the combinatoric selection of the indices in a GQSM index vector and of general IM schemes as described in Section II-A, no duplicate indices may exist within a single index vector.
In light of the above, two sources of such erroneous duplicate convergence are identified: a) high transmit symbol similarity, and b) lack of joint/dependent distribution information between the unit vector variables. In order to mitigate the error from the above identified sources, three modifications to the GQSM transmitter and the UVD-GaBP decoder are proposed in the following.
@@ V-A IQ-Orthogonal Constellation Design
Let us assume that multiple transmit symbols have high similarity, i.e., $|s^{\mathrm{R}}_{p}-s^{\mathrm{R}}_{p^{\prime}}|^{2}\!\approx\!0$. Such behavior is illustrated by observing Fig. 4b and considering $s^{\mathrm{R}}_{p}\approx s^{\mathrm{R}}_{p^{\prime}}$, where the initialized effective symbols $s^{\mathrm{R}}_{p}\cdot\hat{\mathbf{e}}^{\mathrm{R}}_{k_{p}}$ and $s^{\mathrm{R}}_{p^{\prime}}\cdot\hat{\mathbf{e}}^{\mathrm{R}}_{k_{p^{\prime}}}$ are consequently very similar in value. Therefore, the effective symbols appearing in the core MP steps of the UVD-GaBP decoder in equations (28)-(35), may converge to either of the true symbols $s^{\mathrm{R}}_{p}\cdot{\mathbf{e}}^{\mathrm{R}}_{k_{p}}$ or $s^{\mathrm{R}}_{p^{\prime}}\cdot{\mathbf{e}}^{\mathrm{R}}_{k_{p^{\prime}}}$ and potentially produce an erroneous duplicate estimate.
In order to avoid such duplicate symbols in the IQ domain, we propose a simple rotation scheme to $\mathcal{S}^{\mathrm{c}}$, which leads to non-duplicate PAM symbols in both IQ domains. Furthermore, an optimal rotation angle $\theta^{*}$ is obtained to maximize the minimum Euclidean distance between the symbols of the PAM constellation. There exist an optimal rotation angle $\theta^{*}$ for each of the quadrants, but we only consider the solution within the first quadrant, i.e., $\theta\in(0,\frac{\pi}{2})$, which are shown in Table I.
The optimization problem is described by
$D_{\mathrm{min}\!}(\mathcal{S})\triangleq\!\!\underset{s_{m\vphantom{{}^{\prime}}}\!,s_{m^{\prime}}\in\mathcal{S}}{\mathrm{min}}\,\sqrt{|s_{m\vphantom{{}^{\prime}}}\!-s_{m^{\prime}}|^{2}},~{}\forall m\neq m^{\prime}.\vspace{-0.5ex}$ (49)
$M$ 4 16 32 64 128 256
$\theta^{*}$ 0.464 0.245 0.165 0.124 0.082 0.062
@@ V-B Conditional PDF-based Denoiser
$\check{p}^{\mathrm{R}}_{n},\check{t}^{\mathrm{R}}_{n}=\underset{p,t}{\mathrm{argmax}}\;\mathbb{P}(b_{p:n}^{\mathrm{R}}|\mathbf{e}_{t}),\vspace{-1ex}$ (50)
@@
The conditional PMFs does not exhibit a closed-form expression unlike equation (23) for the independent PMFs, and therefore must be obtained via numerical evaluation. However, as the conditional PMFs are fixed for a given codebook, i.e., system parameters $N_{T}$ and $P$, all values can be pre-computed.
@@ V-C Greedy Successive Interference Cancellation (IC)
$\check{p}^{\mathrm{R}},\check{t}^{\mathrm{R}}=\underset{p,t}{\mathrm{argmax}}\;\mathbb{P}(b_{p}^{\mathrm{R}}|\mathbf{e}_{t}),\vspace{-0.75ex}$ (52)
@@
In hand of a converged index of the most confident unit vector variable, IC can be performed on the received signals corresponding to the selected indices as
$\check{y}^{\mathrm{R}}_{n}\triangleq y_{n}-(\mathbf{h}_{n}^{\mathrm{R}})^{\!\mathsf{T}}(s^{\mathrm{R}}_{\check{p}^{\mathrm{R}}_{n}}\!\cdot\!\mathbf{e}_{\check{t}^{\mathrm{R}}_{n}}),\vspace{-0.75ex}$ (53)
and the corresponding variables are nullified from the respective soft-replica, to be repeated for $P$ successive IC iterations.
@@
$\tilde{\mathbf{k}}^{\mathrm{R}}=\mathrm{sort}([{\check{t}^{\mathrm{R}}_{[1]}},\cdots,{\check{t}^{\mathrm{R}}_{[P]}}]),\vspace{-1ex}$ (54)
@@
The complete procedure, incorporating the proposed receiver-side enhancements of Sections V-C and V-B, is provided as a pseudocode in Algorithm 2, and as a schematic diagram in Figure 6. Notice that for $P=1$, the proposed enhancement is not necessary as there is only a single unit vector to evaluate, and the conditional PMF-based denoiser is not applicable as there is no other unit vector to condition the extrinsic beliefs. In addition, the motivation of such enhancements is not necessary for $P=1$, as the duplicate convergence error cannot occur with a single unit vector variable.
@@ VI-A BER Performance
To analyze the performance of the proposed UVD-GaBP algorithm and the enhancement modules in Sec. V, we perform simulations for different system sizes characterized by $N_{T}$ transmit antennas, $N_{R}$ receive antennas, and $P$ pilot symbols from the $M$-ary constellation, including truly massive systems with large $N_{T}$ and $N_{R}$. Specifically, three algorithmic cases as summarized in Table II, are selected to capture and highlight the improvements introduced by each of the three proposed enhancement modules, namely: Algorithm 1 as the baseline UVD-GaBP decoder; Algorithm 2 with conditional denoiser but without successive IC; and finally the full Algorithm 2 with both the conditional denoiser and successive IC. Note that the optimized constellations are employed in all cases considered.
Our first set of results, depicted in Figure 7, compares the bit error rate (BER) performances of the proposed UVD-GaBP decoder for all cases in Table II and the corresponding MFB, in MIMO systems of size $32\times 32$ employing various numbers of symbols $P$.
@@
The BER plots against the energy-per-bit-to-noise-power-spectral-density ratio ($E_{b}/N_{0}$) capture the effects of the duplicate convergence of unit vectors as identified in Section V, highlighting it as a source of errors, which of course does not occur for $P=1$ but can be substantial for $P>1$. The results show, however, that this error is successfully mitigated by the improvements described in Sections V-B and V-C.
In addition, the plots also exhibit error-floors for $P>1$, which still remain while significantly lowered by the proposed enhancement modules. Such error-floor is a convergent phenomenon of MP algorithms in non-ideal high signal-to-noise ratio (SNR) regimes [69, 70], also enlarged when the number of observation nodes is small, i.e., $N_{R}<N_{T}$, whose mitigation has been extensively studied via various methods, including adaptive MP updates and post-processing [71, 72]. However, such extra modifications to further reduce the error-floors are not specific to the IM structure and its detection, and therefore have been considered out of scope of this article.
@@
Furthermore, it is found that sufficiently large MIMO system sizes can sustain larger $P$ values with significantly reduced error-floors. In order to demonstrate this, we offer our next set of results, given in Figure 8, which show the performances achieved by Algorithm 2 in truly massive system scales of $64\times 64$ and $96\times 96$. We remark that simulation results for GQSM systems of such sizes have never been shown before, precisely due to the prohibitive complexity even with the SotA GQSM detection methods [30, 42].
Notice that Figure 8 also compares the performance of the proposed piloted GQSM scheme against a conventional multiplexed (MUX) MIMO scheme, which in order to enable a fair comparison between the two (proposed and conventional) approaches, the total power and transmission rate (in bits/s/Hz) of both systems have been fixed to identical values, and employ the conventional linear GaBP [67, 71] in the detection of the MUX scheme. It can be seen that, piloted GQSM in massive systems, enabled by the proposed UVD-GaBP detector, outperforms a conventional MIMO-MUX system under equal power and resource constraints, without pilots.
In light of the latter finding, one highly promising application and motivation towards the proposed massive IM framework can be highlighted, which is the efficient incorporation of ISAC via the exploitation of the pilot symbols. Specifically, as the IM framework only embeds information in the index domain and not in the symbols themselves, the sensing functionality can be performed with the known (pilot) symbols via sensing methods available in the literature [60, 73, 64, 61, 62, 17] without interference from unknown communication symbols, which is known to be a significant challenge in ISAC systems in which decoding errors made on communication symbol estimates may deteriorate sensing performance [74, 75], while channel estimation errors made over pilots may impact on BER performance [76].